# Modeling Visual Memorability Assessment with Autoencoders Reveals Characteristics of Memorable Images

### v1

@@ 4 Discussion
We observed a strong positive correlation between the model reconstruction error of images and image memorability scores, supporting the hypothesis that more memorable images are inherently more challenging for the autoencoder to reconstruct. Further, our results showed that the distinctiveness of images, as captured by the autoencoder’s latent space, also correlates with memorability. This strong positive correlation indicates that images with more pronounced features, represented by higher normalized distances from their nearest neighbors, are more memorable. We further observed that the correlation between reconstruction error and memorability diminished as the model was exposed to the images repeatedly. This result suggests that the novelty or distinctiveness of an image, which contributes to its memorability, decreases with repeated exposure. This finding generates a new testable hypothesis: repeated exposure reduces an image’s memorability by diminishing its distinctiveness in the latent space. Future experiments can explore how varying levels of exposure affect the underlying features that contribute to memorability, providing deeper insights into the relationship between exposure, distinctiveness, and memory retention. Our approach supports previous studies that have suggested items that stand out from their context are better remembered [4, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40].
@@
Our findings on the correlation of reconstruction errors and image memorability are closely related to the work by Lin and colleagues (2023). They introduced a sparse coding model that used reconstruction error and distinctiveness as metrics for memory formation and retention, focusing on general perceptual memory processes rather than image memorability specifically [15]. Our work differs from theirs in that we specifically target image memorability, demonstrating that reconstruction error can serve as a proxy for memorability, even when reconstructing images themselves in unsupervised settings. Whereas their model focused on compressing image feature embeddings extracted from layers of an object recognition model to explore the broader interplay between perception and memory, our work diverges by directly quantifying visual memorability as an intrinsic attribute of an image. Additionally, while they employed sparse coding, our use of autoencoders, particularly VGG16-based architectures, offers more granular and interpretable representations of image features. This approach allows for clearer insights into which aspects of an image contribute most to memorability, providing a stronger and more interpretable framework for predicting memorability. This makes our model more applicable to practical memorability prediction tasks and provides a stronger framework for interpretability in terms of the underlying image features that influence human memory.
@@
The positive correlation between the autoencoder reconstruction error and memorability score was consistently observed across the different loss functions we tested. Notably, the highest correlation was achieved when perceptual-based loss functions were used to train the model (see Table 1). These results underscore the critical role of selecting appropriate loss functions when modeling human cognitive processes. Specifically, the superior performance of perceptual loss functions, such as LPIPS (SqueezeNet), indicates that perceptual similarity is more important than pixel-level similarity when modeling memorability assessment. The strong correlation between reconstruction error and memorability when LPIPS was used for training highlights the significance of perceptual factors in predicting memorability. This finding has broader implications for designing machine learning systems that aim to replicate human cognitive processes, particularly in domains like visual memory and perception.
@@
The high accuracy of the prediction model (GRU) on the autoencoder’s latent representations in predicting memorability further demonstrated that the compressed representations generated by the autoencoder contain information to predict memorability scores, suggesting the strength of latent space representations as a proxy for human memory assessment. The ability of the model to capture essential features for memorability suggests that latent space representations encode vital information that correlates strongly with human memory recognition.
@@
In addition, the use of Integrated Gradients (IG) for model interpretation provided new insights into the visual features most influential in making an image memorable, addressing a key limitation in earlier works that lacked model interpretability. For instance, visual features such as strong contrasts, distinctive objects, prominent foreground elements, unique textures, and the presence of human figures or actions emerged as key indicators of memorability. Our findings are supported by several established studies in the field. For example, Isola et al. found that semantic content such as the presence of objects and scenes plays a significant role in predicting memorability, with certain types of content (e.g., human figures, interiors) being more memorable than others (e.g., natural scenes, backgrounds) [1]. Kholsa et al. mention that memorable images tend to include distinct objects, people, and salient actions or events, while forgettable images are often characterized by natural landscapes or scenes that lack distinguishable features [5]. Bylinskii et al. also noted that overly simple or uniform images tend to be less memorable [4].
4.1 Limitations and future directions:
@@
While our findings are significant, there are certain limitations to our approach. First, the MemCat dataset, though diverse, may not fully capture the full range of image types and contexts that are relevant to memorability. Future studies should explore broader and more diverse datasets to validate and extend these findings. Second, the use of a VGG16-based autoencoder, though effective, may not be optimal for all types of images. Brain-inspired architectures or other deep learning models should be explored in future research to enhance the model’s capacity to capture different visual features. Third, while Integrated Gradients provided valuable insights into model interpretation, the choice of baseline images and the complexity of the model could influence the reliability of these interpretations. Future research should investigate additional explainable AI techniques to provide a more comprehensive understanding of how visual features influence memorability predictions.
@@
5 Conclusions
Our study presents a computational model for modeling and assessing image memorability, moving beyond traditional human-centric evaluations. By leveraging the VGG16-based autoencoder architecture, which is a deep convolutional neural network (CNN), pre-trained on large-scale image data, we were able to effectively capture high-level visual features contributing to memorability. By investigating reconstruction errors and analyzing latent space representations, we introduce a computational method for understanding the factors that make images memorable. The positive correlation between memorability scores, autoencoder’s reconstruction errors, and latent space distinctiveness highlights the inherent complexity and uniqueness of memorable images. These attributes, which challenge autoencoders, seem to be the same features that make images memorable to human observers.
Modeling memorability assessment using an autoencoder based on deep convolutional neural networks, pre-trained on generalized knowledge from a large dataset, we have demonstrated that deep learning models can effectively capture the key features contributing to human memorability. This research highlights the complex relationship between human cognition and AI capabilities, showing that deep learning models can simulate human memory processes.
Our promising results suggest that this approach could open new avenues for research in both cognitive science and computer vision. This framework offers valuable insights into how artificial intelligence can be used to computationally identify and optimize memorable visual content.
[…]
Supplementary information
The code and scripts required to reproduce the results presented in this paper are available in the following GitHub repository: https://github.com/Elham-Bagheri1/image_memorability.
Declarations
  • Funding: This work is supported by an NSERC Discovery Grant to Y.M. and a Vector Institute Postdoctoral Fellowship to E.B.
  • Conflict of interest/Competing interests: The authors declare no conflict of interest.
  • Ethics approval and consent to participate: This study uses publicly available data, which does not contain identifiable personal data and therefore did not require additional ethical approval.
  • Consent for publication: All authors consent to the publication of this manuscript.
  • Data availability: The data used in this study is publicly available and was not collected by the authors.
  • Materials availability: Not applicable
  • Code availability: The code used for this study is available.
  • Author contribution: E.B. and Y.M. contributed to the conceptualization and design of the study. E.B. was primarily responsible for model development, data analysis, and drafting the manuscript. Y.M. provided supervision, methodological guidance, and critical revision of the manuscript. Both authors reviewed and approved the final version of the manuscript.

### v2

@@ Discussion
We further observed that the correlation between reconstruction error and memorability diminished as the model was exposed to the images repeatedly. This result suggests that the novelty or distinctiveness of an image, which contributes to its memorability, decreases with repeated exposure. This aligns with theories in cognitive psychology indicating that repeated exposure leads to habituation, reducing an image’s perceptual saliency over time. This finding generates a new testable hypothesis: repeated exposure reduces an image’s memorability by diminishing its distinctiveness in the latent space. Future experiments can explore how varying levels of exposure affect the underlying features that contribute to memorability, providing deeper insights into the relationship between exposure, distinctiveness, and memory retention.
Our approach supports previous studies that have suggested items that stand out from their context are better remembered [4, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. However, unlike these earlier studies, our model enables an unsupervised, computational evaluation of memorability, moving beyond human-centric assessment methods and allowing large-scale, automated analysis of visual memory encoding.
@@
Our findings on the correlation of reconstruction errors and image memorability are closely related to the work by Lin et al. [15]. They introduced a sparse coding model that used reconstruction error and distinctiveness as metrics for memory formation and retention, focusing on general perceptual memory processes rather than image memorability specifically. While both studies identify reconstruction error as an important predictor of memorability, our approach differs in several ways. First, we apply autoencoders rather than sparse coding, allowing us to reconstruct images directly while analyzing their latent space representations. Second, our work focuses explicitly on image memorability rather than broader perceptual memory encoding, demonstrating that autoencoder reconstruction error can serve as a direct proxy for memorability. Finally, whereas Lin et al. focused on compressing feature embeddings extracted from an object recognition model, our model directly quantifies memorability as an intrinsic image property.
@@
The positive correlation between the autoencoder reconstruction error and memorability score was consistently observed across the different loss functions we tested. However, the highest correlation was achieved when perceptual-based loss functions were used to train the model. This finding shows the importance of selecting appropriate loss functions when modeling human cognitive processes. This result also suggests that future modeling approaches aiming to capture human cognitive patterns should prioritize perceptual alignment rather than pixel-level fidelity, especially in domains like visual memory and perception.
@@
The reasonable accuracy of the prediction model on the autoencoder’s latent representations in predicting memorability further demonstrated that the compressed representations generated by the autoencoder contain information to predict memorability scores, suggesting the strength of latent space representations as a proxy for human memory assessment. The ability of the model to capture essential features for memorability suggests that latent space representations encode vital information that correlates strongly with human memory recognition.
In addition, the use of Integrated Gradients (IG) for model interpretation provided new insights into the visual features most influential in making an image memorable, addressing a key limitation in earlier works that lacked model interpretability. By directly attributing memorability predictions to specific visual features, our results provide a more concrete mapping between computational representations and human perception.
@@
Our findings are supported by several established studies in the field. For example, Isola et al. found that semantic content such as the presence of objects and scenes plays a significant role in predicting memorability, with certain types of content (e.g., human figures, interiors) being more memorable than others (e.g., natural scenes, backgrounds) [1]. Kholsa et al. mention that memorable images tend to include distinct objects, people, and salient actions or events, while forgettable images are often characterized by natural landscapes or scenes that lack distinguishable features [5]. Bylinskii et al. also noted that overly simple or uniform images tend to be less memorable [4].
An exception to the observed trends was found in the food category, where the correlation between memorability and both reconstruction error and distinctiveness was lower than in other categories. Unlike objects, landscapes, or human figures, food images exhibit greater visual homogeneity in color, texture, and composition, which likely reduces their distinguishability in the model’s latent space. This suggests that food memorability may depend more on semantic, cultural, or personal associations rather than purely visual distinctiveness. Future work should explore incorporating multimodal embeddings (e.g., text-based descriptions, prior exposure modeling) to better capture non-visual memorability cues in such categories.
@@
Limitations and future directions: While our findings are significant, there are certain limitations to our approach. First, the MemCat dataset, though diverse, may not fully capture the full range of image types and contexts that are relevant to memorability. Future studies should explore broader and more diverse datasets to validate and extend these findings. Second, the use of a CNN-based autoencoder, though effective, may not be optimal for all types of images. Brain-inspired architectures or other deep learning models should be explored in future research to enhance the model’s capacity to capture different visual features. Third, while Integrated Gradients provided valuable insights into model interpretation, the choice of baseline images and the complexity of the model could influence the reliability of these interpretations. Future research should investigate additional explainable AI techniques to provide a more comprehensive understanding of how visual features influence memorability predictions.
@@
Our findings have direct applications in AI-driven content generation, advertising, and educational media, where optimizing visual elements for memorability can enhance engagement and information retention. In digital marketing, AI systems could generate ad creatives that maximize recall, while in education, instructors could design learning materials that leverage memorability cues for improved knowledge retention. In medical imaging, this approach could be used to highlight diagnostically important regions by optimizing the memorability of critical visual patterns.
By integrating cognitive science principles with AI-based feature extraction, this study contributes to both theoretical understanding and practical applications in AI-driven memorability modeling. These findings pave the way for novel computational tools that evaluate and enhance visual content based on memorability, ultimately improving user engagement, information retention, and human-computer interaction strategies.
Data Availability
The datasets used in this study are publicly available. MemCat can be downloaded from the MemCat dataset website, and ImageNet is available at the ImageNet download page.
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant to Y.M. and a Vector Institute Postdoctoral Fellowship to E.B.
Author contributions statement
E.B. and Y.M. contributed to the conceptualization and design of the study. E.B. was primarily responsible for model development, data analysis, and drafting the manuscript. Y.M. provided supervision, methodological guidance, and critical revision of the manuscript. Both authors reviewed and approved the final version of the manuscript.
Competing interests
The authors declare that they have no conflicts of interest related to this work.